Hi there, I would appreciate some help with the next problem:
Kristal can wax a floor in 16 minutes. One day her friend Perry helped her and it only took 5.76 minutes. How long would it take Perry to do it alone?

Found 2 solutions by ewatrrr, htmentor:
Answer by ewatrrr(24785)


Hi
Let x represent the time it would take Perry to do it solo
Question states*** Per minute the equalizer
1/16 + 1/x = 1/5.76
Solving for x
36x + 576 = 100x |multiplying thru by 576x so as all denominators= 1
576 = 64x
x = 9 minutes, time it would take Perry to do it alone

Answer by htmentor(1343)

Kristal's rate = 1 floor in 16 min. = 1/16 floors per min.
Working together their rate = 1 floor in 5.76 min. = 1/5.76 floors per min.
We need to determine Perry's rate, so we need to equate their combined rate with the sum of the individual rates.
Let Perry's time = x
In equation form, we can write:
1/16 + 1/x = 1/5.76
Solve for x:
(x+16)/16x = 1/5.76 -> 5.76(x+16) = 16x -> 10.24x = 92.16 -> x = 9
So, Perry would take 9 mins. to do the floor alone.